OpenStudy (anonymous):

place 20 pennies on the table in a row. Now replace every fourth coin wit a nickel. Now replace every third coin with a dime. Now repace every sixth coin with a quarter. What is the value of the 20 coins now on the table?

OpenStudy (anonymous):

Help someone

OpenStudy (anonymous):

if it was me, i would get some change and do it

OpenStudy (anonymous):

i will try and see what i get

OpenStudy (anonymous):

I got $1.35

OpenStudy (anonymous):

i didn't

OpenStudy (anonymous):

i got $1.53

OpenStudy (anonymous):

my coins looked like p p d n p q d n p p d q p p d n p q d n

OpenStudy (anonymous):

8 pennies, 4 nickels, 5 dimes, 3 quarters

OpenStudy (anonymous):

try again, because i do not guarantee that i did not make a mistake

OpenStudy (anonymous):

20/4 = 5 5 of the pennies will be replaced by nickles 20/3 = 6.6 repeating, round down to 6 3 and 4 share the multiple 12, and 18, so two of your nickles are replaced with dimes you now have 11 pennies, 3 nickles, and 6 dimes 20/6 = 3.3 repeating, round down to 3 3 and 6 share the multiples 6, 12, and 18 12 is a dime, and 18 is a dime you have 11 pennies, 2 nickles, 4 dimes, and 3 quarters \[(11*1) + (2*5) + (4+10) + (3*25) = $1.36\]

OpenStudy (anonymous):

how can you have repeating decimal money?

OpenStudy (anonymous):

you don't, you have uneven division which you round off; that's why I rounded all of the numbers down rather than to the nearest digit.

OpenStudy (anonymous):

I rounded the number of coins, not the amount of money. :P

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how can you have a fractional number of coins?

OpenStudy (anonymous):

you replace every 6th coin with a quarter 20/6 = 3.333 coins replaced you can only add 3 coins, because the fourth would land on 24, so you round it off.

OpenStudy (anonymous):

i just got out some change and did it again i got $1.30 this time once more

OpenStudy (anonymous):

I'm gonna run the math one more time, 1 sec

OpenStudy (anonymous):

finally! i got $1.35 just like @Mindless143

OpenStudy (anonymous):

coins # 4,8,16,20 are nickles coins # 3,9,15 are dimes coins # 6,12,18 are quarters 10 pennies, 4 nickles, 3 dimes, 3 quarters \[10 + (4*5)+(3*10)+(3*25) = $1.35\] so $1.35 was correct, I missed a coin or two apparently.